x^3 - 3x^2 + 3x - 9 = 0
Answer by stanbon(48558)

x^2(x-3)+3(x-3)=0
(x-3)(x^2+3)=0
x=3 if your answers are limited to the Real numbers.
If complex are included, then these are also answers:
x=isqrt3 or x=-isqrt3
Cheers,
Stan H.